Circular No. 9058

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 O3 (HILL) R. E. Hill reports his discovery of a comet with the Catalina 0.68-

m Schmidt telescope (discovery observation tabulated below), noting that co-added 30-s exposures show a nuclear condensation elongated northeastsouthwest and a 6′′ × 10′′ coma with a broad, diffuse tail ∼ 2′ long in p.a. 260^{o} (though on July 30.4 UT, the coma < 2′′). After posting on the ‘NEOCP’ webpage, other CCD astrometrists have noted the cometary appearance. Below are given measured coma diameters and tail information from images obtained July 30.0–30.5: 10′′ coma (diffuse with central condensation), 30′′ tail in p.a. 250^{o} (P. Birtwhistle, Great Shefford, Berkshire, England, 0.40-m reflector); strong central condensation with a tail at least 40′′ long in p.a. 255^{o} (L. Buzzi, Varese, Italy, 0.60-m reflector); tail > 1′ long toward p.a. 245^{o} (E. Prosperi, Larciano, Italy, 0.35-m reflector; fifty co-added 30-s exposures taken by P. Camilleri, E. Guido, G. Sostero, and himself); compact nuclear condensation and a broad, diffuse tail 27′′.7 long in p.a. 247^{o}.5 (R. Holmes and H. Devore, Charleston, IL, U.S.A., 0.61-m astrograph); ∼ 5′′ compact coma and a 30′′ tail in p.a. 225^{o} (E. Guido and G. Sostero, remotely using a 0.25-m reflector near Mayhill, NM, U.S.A.; six co-added unfiltered 120-s exposures); round 10′′ coma (J. E. McGaha, Tucson, AZ, U.S.A., 0.36-m reflector; five stacked 60-s exposures); faint tail in p.a. ∼ 225^{o} (W. H. Ryan, Magdalena Ridge 2.4-m reflector).

2009 UT α_{2000} δ_{2000} Mag. Observer
July 29.43294 1^{h}31^{m}27.^{s}03 +10^{o}16′52′′.4 17.5 Hill

The available astrometry, the following preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-O54.

T = 2009 Apr. 21.770 TT  ω = 130.^{o}389
Ω = 180.985 }2000.0
q = 1.88365 AU  i = 12.750

COMET C/2008 Y18 (SOHO) Further to IAUC 9042, another Kreutz sungrazer has been found on

SOHO website images, being extremely faint (mag ∼ 8.5–9) and slightly diffuse in C2 images, and diffuse (mag ∼ 11.5) in STEREO HI1-A images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 Y18 Dec. 25.829 18^{h}24.^{m}5 −24^{o}47′ C2* MK 2009-M41
2009 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green